How much will it cost to purchase materials for a fence if 79 feet are needed and the fencing costs $ 16.63 /yard?

79 feet x ...../..... x ......./....... = $ ...........


1
Expert's answer
2021-01-25T04:20:35-0500

We know,

1 foot = (1/3)yard

So, 79 feet = (79/3) yard

The fence costs, 16.63 $/yard

So, total cost =

79 x 16.63 /3 = $ 437.923

Hence, it costs $ 437.923 to purchase